OpenStudy (anonymous):

Can someone explain how to do this?? Write y=(-5/8)x+3 in standard form using integers.

OpenStudy (anonymous):

Anyone??

OpenStudy (dayakar):

\[y= \frac{ -5 }{ 8 }x +3\] multiply both sides with 8

OpenStudy (dayakar):

can u do

OpenStudy (dayakar):

8y = 8* (-5/8)x +8*3 8y = -5x+24 add both sides 5x

OpenStudy (dayakar):

5x+8y = -5x+5x+24 5x+8y=24

OpenStudy (dayakar):

this is standard form